OpenStudy (anonymous):

a long straight wire carries a current I= 20.0 A flowing vertically upward . If the horizontal component of the earth magnetic field is 2.0 x 10^-5 T . Calculate the resultant magnetic field at a point (a) 10 cm to the east (b) 10 cm to the west ans : a. 6 x 10^-5 to north b. 2 x 10^-5 to south can somebody explain to me how to calculate this question. thanx :)

OpenStudy (anonymous):

 hint ..

OpenStudy (anonymous):

sashagenius,you can calculate very easily if you know two points that (1) the direction of horizontal magnetic field. (2)and the value of magnetic field and direction due to infinite long straight wire at point which is d distance apart from the wire. and using these two concepts calculate the value of magnetic field.. (i) at the right side of the wire the magnetic field due to wire at distance 10cm from it is in same direction of horizontal component of earth magnetic field..i.e B=10^-7*2I/d+Bh...here Bh is horizontal copmponent of earth magnetic field and I is current through wire and d is distance of point from wire. B=10^-7*2*20/.1+2*10^-5=6*10^-5T. (2)in this case these two are in opposite direction simply calculate to take the sign of Bh is -ve.i.e B=10^-7*2I/d-Bh=4*10^-5-2*10^-5=2*10^-5T....here in such a question you must apply vector applications for answering these questions very easily..